OpenStudy (anonymous):

Determine whether the following two lines are parallel, perpendicular, neither, or both. L1: passes through (5, 7) and (2, 8) L2: passes through (0, 5) and (3, 4)

OpenStudy (anonymous):

you need HOW or need ANSWER?

OpenStudy (anonymous):

How would be nice so i can do it without help in the future

OpenStudy (anonymous):

ok, you should find slope of lines that through from points, can you find it?

OpenStudy (anonymous):

i believe it is perpendicular but im not certain

OpenStudy (anonymous):

\[\frac{\Delta y}{\Delta x}\]for L1, it's\[\frac{8-7}{2-5}=-\frac{1}{3}\]and for L2 is\[\frac{4-5}{3-0}=-\frac{1}{3}\]so slopes are equal

OpenStudy (anonymous):

so its parallel?

OpenStudy (anonymous):

yup ;)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw ;)